OpenStudy (anonymous):

There are 10 marbles in a bag, and the marbles are either red or blue. Eric will randomly choose two marbles from the bag, without replacing the first one. If the probability of both marbles' being red is 2 15 , how many RED marbles are in the bag?

OpenStudy (bahrom7893):

(x/10)*(x/9)=2/15

OpenStudy (bahrom7893):

x^2/90=2/15

OpenStudy (bahrom7893):

hmm wait: (x/10)*((x-1)/9)=2/15

OpenStudy (bahrom7893):

there we go that looks better. We remove one red, and one marble from the bag.

OpenStudy (bahrom7893):

(x^2-x)/90=2/15 x^2-x=12 solve for x

OpenStudy (anonymous):

x=4?

OpenStudy (bahrom7893):

yup